# Bootstrapping Language Models with DPO Implicit Rewards

### v1

@@ 4.3 Experience Replay
Though DICE enables us to learn from the response of the current policy, we know that the implicit reward model from the DPO training is not a perfect proxy for human preferences. Solely relying on the implicit reward model may result in forgetting the knowledge inbuilt in the initial policy at the first DPO stage. Inspired by the technique of experience replay [18] in continual learning for preventing catastrophic forgetting and making a good balance between old and new data, we propose to use a mixture of the generated data and the offline preference dataset. While offline preference data is considered to be high-quality, it is off-policy samples; the generated data is closer to the on-policy samples, but the imperfect implicit reward model may introduce noise in labeling the preference. Combining the two can make for a good balance.
Our complete algorithm is summarized in Algorithm 1. During each iteration $t$, we generate $K$ responses ${\mathbf{y}}_{1},{\mathbf{y}}_{2},\ldots,{\mathbf{y}}_{K}$ from the policy $\pi_{\theta^{(t-1)}}$ for each prompt ${\mathbf{x}}$ (3). The preference dataset ${\mathcal{D}}(\alpha)$ is constructed based on the LR rewards $r_{\mathrm{LR}}$, where the response with the highest LR reward is labeled ${\mathbf{y}}_{w}$ and the one with the lowest reward is labeled ${\mathbf{y}}_{l}$. In this process, $\pi_{\theta^{(t-1)}}$ serves as the target policy, and $\pi_{\mathrm{ref}}^{(t-1)}$ serves as the reference policy. At 4, we construct the debiased dataset ${\mathcal{D}}(\alpha^{\star})$ by minimizing the average absolute difference in response length, referring to Eq. 9. Subsequently, a mixed dataset ${\mathcal{D}}_{t}$ is created by sampling $\gamma$ proportion of data from ${\mathcal{D}}_{\mathrm{offline}}$ and $(1-\gamma)$ proportion of data from ${\mathcal{D}}(\alpha^{\star})$ (5). DPO training is then conducted on ${\mathcal{D}}_{t}$ using $\pi_{\theta^{(t-1)}}$ as both the initial policy and the reference policy, resulting in the updated policy $\pi_{\theta^{(t)}}$ (6). Finally, at 7, we set $\pi_{\theta^{(t-1)}}$ as the new reference policy $\pi^{(t)}_{\mathrm{ref}}$.
@@ 5.1 Experiment Setup
Base Models and Datasets. In this paper, we adopt Llama-3-8B-DPO [^2] https://huggingface.co/princeton-nlp/Llama-3-Base-8B-SFT-DPO and zephyr-7B-beta [^3] https://huggingface.co/HuggingFaceH4/zephyr-7b-beta as our base models. Both models are trained following the pipeline of Zephyr [26]. Llama-3-8B-DPO is trained based on meta-llama/Meta-Llama-3-8B [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B, developed by Meng et al. [13]. zephyr-7B-beta is fine-tuned based on mistralai/Mistral-7B-v0.1555https://huggingface.co/mistralai/Mistral-7B-v0.1. We adopt UltraFeedback preference dataset [4] and randomly sample a subset of around 10k preference pairs as the offline dataset ${\mathcal{D}}_{\mathrm{offline}}$. Our experiment aims to show how much the language model can improve from a DPO-ed model and a subset of the preference dataset that was used to conduct the initial DPO training.
Response Generation and Dataset Construction. At the beginning of each round, we first sample responses from the current policy, with temperature $T=0.9$, $p=1.0$ for the Llama3 setting and $T=0.7$, $p=0.9$ for the Zephyr setting. We sample with different random seeds to get $K=16$ diverse responses for each prompt. We then reward each prompt-response pair by the implicit reward model (Eq. 4) and incorporate length-regularized reward shaping (Eq. 8) to get the debiased dataset ${\mathcal{D}}(\alpha^{\star})$ with the optimal regularization strength $\alpha^{\star}$. The final dataset is a mixture of ${\mathcal{D}}(\alpha^{\star})$ and ${\mathcal{D}}_{\mathrm{offline}}$.
@@
Baselines. We evaluate the following baseline methods that are applicable to the setting of this paper:
@@
  • Offline DPO w/ new ref: similar to offline DPO but we assign the current policy as the new reference model, while we use a fixed reference model in Offline DPO. This corresponds to $\gamma=1$.
  • DPO with prompted rewards: similar to Self-rewarding LM [28], where they prompt the LLM itself as a preference judge to construct new preference pairs and iteratively fine-tune the LLM with the DPO algorithm. In their case, the judge capability is learned by supervised fine-tuning on an evaluation fine-tuning dataset. We exploit our base model to perform LLM-as-a-Judge directly as it can follow the judge instructions well. In the experiment, we call it LLM-as-a-Judge. The LLM-as-a-Judge prompt template can be found in Appendix A.
Table 1: AlpacaEval 2.0 evaluation of various models (detailed in Baselines in Section 5.1) in terms of both normal win rates (WR) and length-controlled win rates (LC) in percentage (%).
@@
LC WR Avg. Len. LC WR Avg. Len.
Base 12.69 10.71 1444 18.20 15.50 -
Offline DPO Iter 1 13.40 11.10 1544 20.22 18.33 1750
Offline DPO Iter 2 4.96 5.47 2178 21.04 19.21 1761
Offline DPO (w/ new ref) Iter 1 13.40 11.10 1544 22.29 19.96 1775
Offline DPO (w/ new ref) Iter 2 4.58 5.27 2296 22.50 20.18 1737
LLM-as-a-Judge Iter 1 17.81 15.36 2572 20.30 21.31 2029
LLM-as-a-Judge Iter 2 14.14 17.89 2916 21.80 22.42 2051
DICE Iter 1 19.03 17.67 1848 25.08 25.77 2201
DICE Iter 2 20.71 20.16 1879 27.55 30.99 2427
    We note that the results of Llama-3-8B-DPO base are obtained from Meng et al. [13].
Evaluation. We evaluate our method by AlpacaEval 2.0 [12]. AlpacaEval 2.0 is an LLM-based automatic evaluation benchmark. It employs AlpacaFarm [5] as its prompts set composed of general human instructions. The model responses and the reference responses generated by GPT-4-Turbo are fed into a GPT-4-Turbo-based annotator to be judged. We follow the standard approach and report both the win rate (WR) and the Length-Controlled win rate (LC) [6] over the reference responses.
@@ 5.2 Main Results
Table 2: AlpacaEval 2.0 leaderboard results.
@@
Compare with the models on the leaderboard. Compared with the models on the public leaderboard shown in Table 2, DICE-Llama3 8B performs better than the official instruct version of Llama3 by a non-trivial margin, $4.63\%$. Regarding the closed-source models, it achieves better performance than Gemini Pro with only 8B parameters and does not require any in-house data or external reward model.
Though DICE works best with DPO as it makes the iterative training possible (because the implicit reward model for the next round can be naturally derived using the updated policy), we would like to check if the dataset generated by DICE can also improve the base model with other DAP algorithms. In the Zephyr setting, we tune the policy model using DICE-generated dataset (at the first round) with KTO [7], IPO [1], and Hinge loss proposed in Zhao et al. [29]. The training follows the protocol described in Section 5.1. The results in Table 3 show that all DAP algorithms benefit from the newly generated data by the current policy and DPO implicit reward model with LR reward shaping, demonstrating LC win rates higher than the offline DPO. Meanwhile, DICE works best with DPO while IPO is the runner-up. It should be noted that DICE-IPO has a serious length exploitation issue even on the dataset where LR reward shaping has been applied.
Table 3: DPO rewards are compatible with other direct preference optimization algorithms.
Method LC WR Avg. Len.
Offline DPO 13.40 11.10 1544
DICE-IPO 18.51 19.49 2729
DICE-KTO 14.88 12.16 1541
DICE-Hinge 15.92 15.57 1970
DICE-DPO 19.03 17.67 1848
Table 4: Effects of the Length Regularized Weight.
@@ 5.4 Ablation Study
Figure 3: AlpacaEval 2.0 LC Win Rate across different experience replay ratio ($\gamma$) in the Zephyr setting. The highest LC Win Rate is reported via text.
@@
Effects of experience replay. The experience replay results in a new mixed dataset in which $\gamma$ fraction of the data is from the offline dataset and $(1-\gamma)$ fraction of the data is from the generated dataset. E.g., we use data only from the generated dataset if $\gamma=0.0$. We run experiments in the Zephyr setting with $\gamma\in\{0.0,0.25,0.5,0.75,1.0\}$ and conduct DICE in total of two rounds. From the results shown in Figure 3, we find $\gamma=0.5$ provides the best performance. The results satisfy our expectations. With only offline preference data, the DPO optimizes the current policy with off-policy samples that are further away from its distribution. With only its own generated data, the model may keep reinforcing its current “belief”, potentially leading to catastrophic forgetting. An intermediate value $\gamma=0.5$ finds a good balance. Additional results in the Llama3 setting can be found in Appendix C.
6 Limitations, Future Work and Broader Impacts
@@
Limitations. One of the primary limitations of our work is the reliance on the DPO training prior to bootstrapping. If the implicit reward model is not well-trained, it can lead to a collapse of the training pipeline. This challenge is not unique to our approach; the classic RLHF pipeline is also struggling when its reward model is not well-trained. Another limitation is the lack of continued improvement over many iterations. Similar to other research, such as the work by [28], which enhances the policy model without an external reward model, we did not observe continuous improvement in our model beyond three iterations. This issue highlights an open question within this field regarding the iterative enhancement of policy models.
@@
Future Work. Future research could explore the rewarding capabilities of models trained using other DPO variants, such as KTO and IPO. Investigating whether these variants can offer general rewards similar to those provided by DPO-ed policy would be valuable. Another promising direction involves developing methods that enable continuous improvement of the policy model over iterations without degradation. Additionally, investigating a theoretical understanding of the policy learned through self-bootstrapping could provide deeper insights into the mechanics of our approach and facilitate further advancements.
Broader Impacts. Due to its simplicity in implementation and its ability to enhance the policy model effectively, our method has the potential to be misused, for instance, in generating harmful content. The ease of use and the capability to produce stronger policy models might attract malicious actors who seek to exploit these features for unethical purposes. It is crucial to address these concerns by developing safeguards and ethical guidelines to prevent misuse.

### v2

@@ 3.2 Experience replay
Though DICE enables us to learn from the response of the current policy, we know that the implicit reward model from the DPO training is not a perfect proxy for human preferences. Solely relying on the implicit reward model may result in forgetting the knowledge inbuilt in the initial policy at the first DPO stage. Inspired by the technique of experience replay (Rolnick et al., 2019) in continual learning for preventing catastrophic forgetting and making a good balance between old and new data, we propose to use a mixture of the generated data and the offline preference dataset. While offline preference data is considered to be high-quality, it is off-policy samples; the generated data is closer to the on-policy samples, but the imperfect implicit reward model may introduce noise in labeling the preference. Combining the two can make for a good balance. This idea is also related to deep Q-learning with demonstrations (Hester et al., 2018), which demonstrates that incorporating supervised learning from an offline dataset can accelerate online reinforcement learning.
Our complete algorithm is summarized in Algorithm 1. During each iteration $t$, we generate $K$ responses ${\mathbf{y}}_{1},{\mathbf{y}}_{2},\ldots,{\mathbf{y}}_{K}$ from the policy $\pi_{\theta^{(t-1)}}$ for each prompt ${\mathbf{x}}$ (3). The preference dataset ${\mathcal{D}}(\alpha)$ is constructed based on the LR rewards $r_{\mathrm{LR}}$, where the response with the highest LR reward is labeled ${\mathbf{y}}_{w}$ and the one with the lowest reward is labeled ${\mathbf{y}}_{l}$. In this process, $\pi_{\theta^{(t-1)}}$ serves as the target policy, and $\pi_{\theta^{(t-2)}}$ serves as the reference policy. For $t=1$, we denote the reference policy in the implicit reward model as $\pi_{\theta^{(-1)}}$, corresponding to the reference policy used in the initial DPO training. At 4, we construct the debiased dataset ${\mathcal{D}}(\alpha^{\star})$ by minimizing the average absolute difference in response length, referring to Eq. 6. Subsequently, a mixed dataset ${\mathcal{D}}_{t}$ is created by sampling $\gamma$ proportion of data from ${\mathcal{D}}_{\mathrm{offline}}$ and $(1-\gamma)$ proportion of data from ${\mathcal{D}}(\alpha^{\star})$ (5). Finally, DPO training is conducted on ${\mathcal{D}}_{t}$ using $\pi_{\theta^{(t-1)}}$ as both the initial policy and the reference policy, resulting in the updated policy $\pi_{\theta^{(t)}}$ (6).
@@ 4.1 Experiment setup
Base Models and Datasets. We adopt Llama-3-8B-DPO [^1] https://huggingface.co/princeton-nlp/Llama-3-Base-8B-SFT-DPO and zephyr-7B-beta [^2] https://huggingface.co/HuggingFaceH4/zephyr-7b-beta as our base models. Both models are trained following the pipeline of Zephyr (Tunstall et al., 2023) on the UltraFeedback (Cui et al., 2023) dataset. Llama-3-8B-DPO is trained based on meta-llama/Meta-Llama-3-8B , developed by Meng et al. (2024). zephyr-7B-beta is fine-tuned based on mistralai/Mistral-7B-v0.1 . We randomly sample a subset of around 10k preference pairs from UltraFeedback as the offline dataset ${\mathcal{D}}_{\mathrm{offline}}$ for our fine-tuning experiments. Our experiments aim to show how much the language model can improve from a DPO-tuned model and a subset of the preference dataset that was used to conduct the initial DPO training.
Response Generation and Dataset Construction. At the start of each round, we sample responses from the current policy, with temperature $T=0.9$, $p=1.0$ for the Llama3 setting and $T=0.7$, $p=0.9$ for the Zephyr setting. We sample with different random seeds to get $K=16$ diverse responses for each prompt. We then reward each prompt-response pair by the implicit reward model (Eq. 4) and incorporate length-regularized reward shaping (Eq. 5) to get the debiased dataset ${\mathcal{D}}(\alpha^{\star})$ with the optimal regularization strength $\alpha^{\star}$. The final dataset is a mixture of ${\mathcal{D}}(\alpha^{\star})$ and ${\mathcal{D}}_{\mathrm{offline}}$.
@@
Baselines. We evaluate the following baseline methods applicable to the setting of this paper:
@@
  • Offline DPO w/ new ref: similar to offline DPO but we assign the trained policy after each round as the new reference model for the next round. This corresponds to $\gamma=1$.
  • DPO with prompted rewards: similar to self-rewarding LM (Yuan et al., 2024), where they prompt the LLM itself as a preference judge to construct new preference pairs and iteratively fine-tune the LLM with the DPO algorithm. In their case, the judge capability is learned by supervised fine-tuning on an evaluation fine-tuning dataset. We exploit our base model to perform LLM-as-a-Judge directly as it can follow the judge instructions well. In the experiment, we call it LLM-as-a-Judge. The LLM-as-a-Judge prompt template can be found in Appendix E.
Table 1: Results of AlpacaEval 2 and Arena-Hard across two base models. LC and WR denote length-controlled and raw win rate in percentage (%) respectively.
@@
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
LC WR LC LC WR LC
Base 12.69 10.71 10.17 18.20 15.50 22.32
Offline DPO Iter 1 13.40 11.10 14.13 20.22 18.33 24.78
Offline DPO Iter 2 4.96 5.47 1.89 21.04 19.21 24.04
Offline DPO (w/ new ref) Iter 1 13.40 11.10 13.10 22.29 19.96 22.2
Offline DPO (w/ new ref) Iter 2 4.58 5.27 5.15 22.50 20.18 23.55
LLM-as-a-Judge Iter 1 15.36 17.81 13.96 20.30 21.31 26.46
LLM-as-a-Judge Iter 2 14.14 17.89 13.47 21.80 22.42 27.99
DICE Iter 1 19.03 17.67 18.15 25.08 25.77 33.52
DICE Iter 2 20.71 20.16 18.45 27.55 30.99 39.13
    We note that the results of Llama-3-8B-DPO base are obtained from Meng et al. (2024).
Evaluation. We evaluate our method by AlpacaEval 2 (Li et al., 2023b) and Arena-Hard (Li et al., 2024). Both are LLM-based automatic evaluation benchmarks and have been widely adopted by the community. AlpacaEval 2 employs AlpacaFarm (Dubois et al., 2023) as its prompts set composed of general human instructions. The model responses and the reference responses generated by GPT-4-Turbo are fed into a GPT-4-Turbo-based annotator to be judged. We follow the standard approach and report both the win rate (WR) and the Length-Controlled win rate (LC) (Dubois et al., 2024) over the reference responses. Arena-Hard is a recently released benchmark, incorporating 500 well-defined technical problem-solving queries. Due to the expensive evaluation cost of Arena-Hard, we follow the official guidance and use a strong open-source model mistralai/Mistral-Large-Instruct-2407 (123B) [^3] https://huggingface.co/mistralai/Mistral-Large-Instruct-2407 as the judge model.
@@ 4.2 Main results
Table 2: AlpacaEval 2 leaderboard results.
@@
Comparison with the models on the leaderboard. Compared with the models on the public leaderboard shown in Table 2, DICE-Llama3 8B performs better than the official instruct version of Llama3 by a non-trivial margin, $4.63\%$. Regarding the closed-source models, it achieves better performance than Gemini Pro with only 8B parameters and does not require any in-house data or external reward model.
Table 3: AlpacaEval 2 results showing DPO rewards are compatible with other DAP algorithms.
Method Offline DICE
LC WR LC WR
DPO 13.40 11.10 19.03 17.67
IPO 14.83 16.16 18.51 19.49
KTO 13.92 10.65 14.88 12.16
Hinge 13.51 12.45 15.92 15.57
Table 4: AlpacaEval 2 results showing effects of the Length Regularized Weight.
@@ 4.3 DICE is compatible with other direct alignment from preference Algorithms
Though DICE works best with DPO as it makes the iterative training possible (because the implicit reward model for the next round can be naturally derived using the updated policy), we would like to check if the dataset generated by DICE can also improve the base model with other Direct Alignment from Preference (DAP) algorithms. In the Zephyr setting, we tune the policy model using DICE-generated dataset (at the first round) and the offline dataset with KTO (Ethayarajh et al., 2024), IPO (Azar et al., 2024), and Hinge loss proposed in Zhao et al. (2023). The training follows the protocol described in Section 4.1. The results in Table 3 show that all DAP algorithms benefit from the newly generated data by the current policy and DPO implicit reward model with LR reward shaping, demonstrating LC win rates higher than their offline counterparts. Notably, DICE shows the greatest improvement for DPO.
4.4 Comparing DICE with internal reward model
Our experiment setup assumes access to an offline preference dataset and a DPO-tuned language model. To further improve the language model, one may choose to train a new scalar reward model and then conduct iterative DPO. As this scalar reward model is trained solely on the offline preference dataset without utilizing the external resources, we call it the internal reward model (IntlRM). In this section, we compare the IntlRM with implicit reward model to evaluate their performance in our self-alignment setting.
Base Models and Datasets. We utilize mistral-7b-sft-beta [^4] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta as the base model. We adopt the full UltraFeedback preference dataset (Cui et al., 2023), comprising 60k preference pairs.
Training Details. For the implicit reward, we use zephyr-7B-beta along with its reference model mistral-7b-sft-beta. The scalar reward model is trained using OpenRLHF (Hu et al., 2024) framework, adhering to the recommended training parameters [^5] https://github.com/OpenRLHF/OpenRLHF/blob/main/examples/scripts/train_rm_llama.sh. This setup ensures that both the implicit reward model and IntlRM use the same base model (mistral-7b-sft-beta) and are trained on the same dataset (UltraFeedback).
Evaluation. We evaluate the reward models based on the alignment rate of the model with the preference labels provided by GPT-4o. The alignment rate is defined as $m/n$, where $n$ is the total number of prompt-response pairs and $m$ is the number of labels matching GPT-4o. As our goal is to provide preference labels for the model-generated responses to enable the subsequent DPO training, we sample 500 $({\mathbf{x}},{\mathbf{y}}_{1},{\mathbf{y}}_{2})$ tuples from the preference dataset during the first iteration of the DICE experiment under the Zephyr setting. Different reward models are then queried to provide the preference labels.
Table 5: Alignment rate of the reward model with the preference labels provided by GPT-4o.
Method IntlRM ERM-555k DPO Implicit Reward
Alignment Rate 0.624 0.656 0.698
In the experiment, we include an external reward model, ERM-555k [^6] https://huggingface.co/OpenRLHF/Llama-3-8b-rm-mixture, trained by OpenRLHF maintainers on 555k preference data, as a stronger baseline. In Table 5, we observe that the DPO implicit reward achieves higher accuracy than the IntlRM trained on an equivalent amount of preference data. Furthermore, it surpasses the performance of the ERM-555k trained on substantially more data. This implies that the implicit reward model offers advantages when evaluating its own generated data, though the scalar reward models in general excel on a wider range of tasks when the preference data is abundant, as demonstrated by ArmoRM-Llama3-8B-v0.1, which was trained on 1M preference data. Based on these findings, we argue that the implicit reward is a competitive option in self-alignment settings.
@@ 4.5 Ablation study
Figure 3: AlpacaEval 2 LC Win Rate across different experience replay ratios ($\gamma$) in the Zephyr setting. The highest LC Win Rate is reported via text.
@@
Effects of experience replay. The experience replay results in a new mixed dataset in which $\gamma$ fraction of the data is from the offline dataset and $(1-\gamma)$ fraction of the data is from the generated dataset. E.g., we use data only from the generated dataset if $\gamma=0.0$. We run experiments in the Zephyr setting with $\gamma\in\{0.0,0.25,0.5,0.75,1.0\}$ and conduct DICE in total of two rounds. From the results shown in Figure 3, we find $\gamma=0.5$ provides the best performance. The results satisfy our expectations. With only offline preference data, the DPO optimizes the current policy with off-policy samples that are further away from its distribution. With only its own generated data, the model may keep reinforcing its current “belief”, potentially leading to catastrophic forgetting. An intermediate value $\gamma=0.5$ finds a good balance. Additional results in the Llama3 setting can be found in Appendix D.
5 Limitation and future work
@@
Limitations. One of the primary limitations of our work is the reliance on the DPO training prior to bootstrapping. If the implicit reward model is not well-trained, it can lead to a collapse of the training pipeline. This challenge is not unique to our approach; the classic RLHF pipeline is also struggling when its reward model is not well-trained. Another limitation is the lack of continued improvement over many iterations. Similar to other research, such as the work by Yuan et al. (2024), which enhances the policy model without an external reward model, we did not observe continuous improvement in our model beyond three iterations. This issue highlights an open question within this field regarding the iterative enhancement of policy models.
@@
Future Work. Future research could explore the rewarding capabilities of models trained using other DPO variants, such as KTO and IPO. Investigating whether these variants can offer general rewards similar to those provided by DPO-tuned policy would be valuable. Another promising direction involves developing methods that enable continuous improvement of the policy model over iterations without degradation. Additionally, investigating a theoretical understanding of the policy learned through self-bootstrapping could provide deeper insights into the mechanics of our approach and facilitate further advancements.
6 Related work
Self-Improving Fine-Tuning. Many studies explore fine-tuning language models without a large amount of human annotation (Huang et al., 2022; Li et al., 2023a; Sun et al., 2023; 2024; Yuan et al., 2024; Chen et al., 2024b; Zhang et al., 2025). Starting from an SFT model, Sun et al. (2023) collect the preference labels by prompting the SFT model itself to choose the preferred response based on principles, training a principle-driven reward model, and optimizing via PPO. Yuan et al. (2024) construct a preference dataset by their own supervised fine-tuned model trained on instruction following data and evaluation fine-tuning data, followed by DPO training. Different from these approaches, which align models by leveraging pretrained knowledge with minimal seed data, our work focuses on further improving a DPO-tuned model through bootstrapping with its implicit rewards.
On-Policy Sampling in Preference Tuning. DPO and its variants are widely used for their simplicity, but their offline nature often limits policy learning (Guo et al., 2024). Guo et al. (2024) show that offline DPO quickly overfits the preference dataset, while it performs much better and is more stable if online feedback can be provided to their on-policy samples [^7] On-policy samples: the data collected while following the current policy that is being optimized.. Tajwar et al. (2024) further analyze the properties of different preference tuning methods, concluding that on-policy sampling enhances performance and efficiency. When pure on-policy sampling is infeasible, using data closer to on-policy samples, such as in iterative DPO with external rewards (Tran et al., 2023; Dong et al., 2024; Xiong et al., 2024; Ding et al., 2024; Pang et al., 2024; Muldrew et al., 2024) or internal rewards (Yuan et al., 2024; Furuta et al., 2025), can still be beneficial. Similarly, our approach enables us to train the policy model with the preference data which is closer to on-policy samples than the offline dataset without any external reward models. We hypothesize this is one of the main gain sources of our approach.
DPO Implicit Rewards. Rafailov et al. (2024a) recently studied DPO from a token-level MDP perspective, revealing that DPO-tuned models implicitly parameterize token-wise dense reward functions. Building on similar theoretical fundations, Zhong et al. (2024) pretrains a DPO model to serve as a standalone dense reward generator for PPO training. Unlike these works, we leverage the implicit reward model as an outcome reward model to bootstrap itself for self-alignment. Implicit reward models have also been used for selecting the pairwise data (Yang et al., 2024; Chen et al., 2024a) to improve the annotation efficiency in preference tuning.